In one car rental shop they charge $20 per day with car with an additional 14 cents per mile driven. Another car rental shop charges $22 per day with an additional 10 cents per mile driven. How many miles will the cost of both rental car fees be the same?
Answer by Maths68(1474)

Let
Miles = x
Day = d
Change dollar into cents
$20=20*100=2000cents
$22=22*100=2200cents
2000*d+x*14 = 2200*d+x10
2000d+14x=2200d+10x
14x-10x=2200-2000
4x=200
x=200/4
x=50
For 50 miles the cost of both rental car fees will be the same


Check
======
Miles = 50
Day = 1
One Company will Charge for one day and 50 miles
1 day charges = $20
50 miles charges = 14*50=700cents=$7
Total Charges = 20+7=$27


Other Company will Charge for one day and 50 miles
1 day charges = $22
50 miles charges = 10*50=500cents=$5
Total Charges = 22+5=$27